OpenStudy (anonymous):

Solve each quadratic equation. Show your work. (2x – 1)(x + 7) = 0

OpenStudy (anonymous):

let each factor with equals to zero and solve for x (2x -1) = 0 x= 1/2 do the same with (x+7) Hope it helps

OpenStudy (anonymous):

i'm not sure is that right (x+7)=0 x=7

OpenStudy (anonymous):

No , x = -7

OpenStudy (anonymous):

so x+7=0is that right

OpenStudy (anonymous):